OpenStudy (anonymous):

I'm given that, \[f(x)= e^{1+x}-e^{1-x}\] and that \[g(x) = f^{-1}(x)\] I need [help!] to evaluate \[g'(1)\]

OpenStudy (anonymous):

So the way I tried to work through it was to change f(x) to \[f(x) = e \times e^x- e \times e^{-x}\]then find the inverse of f(x) by changing x and y...but I couldn't isolate for y so I tried using implicit differentiation and everything went down the toilet. I have a solution to the problem but wanted to know if there was an easier way to do it

OpenStudy (anonymous):

First I think you want to solve \(x\) where\[ f(x) = 1 \]

OpenStudy (anonymous):

Ah, why would I do that? ;o

OpenStudy (anonymous):

Hmmm, well, we can start with the derivative I guess.

OpenStudy (anonymous):

\[ g(f(x)) = x \]

OpenStudy (anonymous):

To make the notation less confusing, let \(y=f(x)\). Then \[ g(y) = x \]We can differentiate this with implicit differentiation.

OpenStudy (anonymous):

It gives me the wrong answer...unless the sol'n I'm given is wrong

OpenStudy (anonymous):

differentiating implicitly

OpenStudy (anonymous):

Well, what did you for the derivative?

OpenStudy (anonymous):

\[y=e \times e^{x} - e \times e^{-x}\]\[y=e(e^{x}-\frac{ 1 }{ e^{x}})\]then I switched x and y so that it's the inverse--is that correct to do or is that my mistake? But I'll continue, \[\frac{ x }{ e }=e^{y}-e^{-y}\]\[xe^{-1}=e^{y}-e^{-y}\]\[e^{-1}=e^{y} \times \frac{ dy }{ dx }-e^{-y} \times \frac{ -dy }{ dx }\]\[\frac{ dy }{ dx }=\frac{ e^{-1} }{ e^{y}+e^{-y} }\]

OpenStudy (anonymous):

sorry it takes me so much time to make eqns ;s

OpenStudy (anonymous):

Okay, what did you do after that?

OpenStudy (anonymous):

Nothing, I got confused. I have no value for y, right? I don't know what to dooooo

OpenStudy (anonymous):

Remember that \(y\) is the inverse function.

OpenStudy (anonymous):

You have \(x=1\). And so \(y=g(1)\).

OpenStudy (anonymous):

You asked why I said you should find \[ f(x)=1 \]

OpenStudy (anonymous):

;o so if I solve for g(1) or the x value for which f is 1 and plug it in I'll get the answer? But the answer's insane

OpenStudy (anonymous):

Ok?

OpenStudy (anonymous):

Can't you at least try?

OpenStudy (anonymous):

And how do you know what sort of answer you should get anyway?

OpenStudy (anonymous):

Because someone showed me their solution ^^ I just don't understand it

OpenStudy (anonymous):

I can't solve for when f(x)=1, but I did try taking the ln of both sides isn't helping. Stuck at\[x-1=\ln(e^{2x}-1)\]

OpenStudy (anonymous):

\[ 1 = e(e^x-e^{-x})\implies e^{-1} = e^x-e^{-x} = 2\sinh(x) \]

OpenStudy (anonymous):

loool what's sinh

OpenStudy (anonymous):

\[ \sinh(x) = \frac{e^x-e^{-x}}{2} \]

OpenStudy (anonymous):

I'm not expected to know that lol I give up on this question

OpenStudy (anonymous):

\[ \sinh^{-1}(x) = \ln\left(x+\sqrt{x^2+1}\right) \]

OpenStudy (anonymous):

They never talked about \(\sinh\)? Okay, is the topic of the lesson?

OpenStudy (anonymous):

no they haven't, and no topic it's just a question